3. A 0.649-g sample containing only K2SO4 and (NH4)2so4 was dissolved in water and treated with Ba (NO3)2 to precipitate all so as Baso4. Find the weight percent of K2SO in the sample if 0.977 g of precipitate was formed.
Answer #1

Given,

Mass of K2SO4 + Mass of (NH4)2SO4 = 0.649 g

Mass of precipitate(BaSO4) formed = 0.977 g

Assume,

Mass of K2SO4 = "x"

Thus,

Mass of (NH4)2SO4 = (0.649 - x)

Now, Calculating the number of moles of precipitate(BaSO4) from the given mass,

= 0.977 g BaSO4 x ( 1 mol / 233.38 g)

= 0.004186 mol BaSO4

Now, the separate balanced reactions for K2SO4 and (NH4)2SO4 are,

K2SO4(aq) + Ba(NO3)2\rightarrow BaSO4(s) + 2KNO3(aq)

Similarly,

(NH4)2SO4(aq) + Ba(NO3)2\rightarrow BaSO4(s) + 2NH4NO3(aq)

Now, the moles of K2SO4 and (NH4)2SO4 are,

Moles of K2SO4 = (x / 174.25)

Similarly,

Moles of (NH4)2SO4 = [(0.649-x) / 132.14]

Now, using the moles of  K2SO4 and moles of (NH4)2SO4, calculating the number of moles of precipitate forming from the balanced chemical reaction,

= (x / 174.25) mol K2SO4 x ( 1 mol BaSO4 /1 mol K2SO4)

= (x / 174.25) mol BaSO4

Similarly,

= [(0.649-x) / 132.14] mol (NH4)2SO4 x ( 1 mol BaSO4 /1 mol (NH4)2SO4)

= [(0.649-x) / 132.14] mol BaSO4

Now,

Moles of BaSO4 formed from K2SO4 + Moles of BaSO4 formed from (NH4)2SO4 = 0.004186 mol BaSO4

(x / 174.25) + [(0.649-x) / 132.14] = 0.004186

(132..14x + 113.09 - 174.25x)/ 23025.4 = 0.004186

(132..14x + 113.09 - 174.25x) = 96.39

16.6969 = 42.11x

x = 0.3965

Thus, Mass of K2SO4 in the sample = 0.3965 g

Now, We know, the weight percent formula,

Weight percent of K2SO4 = ( Weight of K2SO4 / Total weight of sample) x 100

Weight percent of K2SO4 = (0.3965 g / 0.649 g) x 100

Weight percent of K2SO4 = 61.1 %